OpenStudy (anonymous):

Decide wheter the equation is an identity, a conditional, or contradiction. 5(2x-31)=10x-155

OpenStudy (anonymous):

identity (meaning always true for all X)

OpenStudy (cwrw238):

can you distribute 5 over the parentheses ?

OpenStudy (cwrw238):

5(2x - 31) = ?

OpenStudy (anonymous):

I don't think so. The questions just asked if this was an identity, a conditional of contradiction.

OpenStudy (cwrw238):

right but doing the above operation will make it clearer about what it is = 5*2x - 5*31 = 10x - 155

OpenStudy (cwrw238):

now you have 10x - 155 = 10x - 155

OpenStudy (cwrw238):

now does that comply with future's definition?